# Everybody Prune Now: Structured Pruning of LLMs with Only Forward Passes

### v3

@@ Abstract
Structured pruning is a promising approach to create smaller, faster LLMs. However, existing methods typically rely on backward passes, which can inflate memory requirements and compute costs. In this work we introduce Bonsai, a gradient-free structured pruning method that eliminates the need for backpropagation, significantly reducing memory requirements and compute costs while achieving state-of-the-art pruning performance. Bonsai uses forward-pass-only perturbative pruning to enable efficient compression of large models on a broader range of hardware configurations. Unlike existing structured pruning approaches, Bonsai not only achieves better compression with fewer resources, but also produces models that are twice as fast as those generated by semi-structured pruning. As a concrete demonstration, we use Bonsai to prune an 8B LLaMA-3 model to 50% sparsity on a single A6000 GPU–a task infeasible with backprop-based methods, which require $2$-3×3$\times$3 × memory. Our results show that removing backprop as a requirement not only enables pruning larger models on constrained hardware but can also lead to state-of-the-art efficiency and performance.
@@ 1 Introduction
Figure 1: Perplexity versus inference speed-up of pruned models for methods that use only forward passes on the parent model. At any given target perplexity, Bonsai produces the fastest model, resulting in improved latency and throughput. Circle sizes $\propto$ model’s memory footprint.
@@
In light of this challenge, we present Bonsai, a structured pruning approach that operates exclusively with forward passes through the parent model. By eliminating the need for backward passes for pruning, Bonsai dramatically reduces memory requirements that enable pruning of models that would otherwise be infeasible on available hardware, and can do so while outperforming existing structured and unstructured pruning approaches (Fig 1). Rather than computing gradients to determine module importance, Bonsai employs an innovative perturbative approach, estimating module importances by evaluating the performance of sub-models using only inference.
We make this perturbative approach tractable through several key innovations. First, we formulate module importance estimation as an underdetermined regression problem, allowing us to infer the importance of a large number of modules by exploring a manageable number of random sub-models. This fundamentally differs from previous pertubative approaches, which require evaluating approximately as many sub-models as there are modules (Ancona et al., 2020), making them intractable for LLMs. Second, we use informative priors derived from unstructured pruning (Han et al., 2015; Sun et al., 2023) to guide sub-model exploration, yielding better estimates of module relevance with fewer evaluations, instead of instantiating sub-models by dropping modules with equal likelihood (Kang et al., 2023). Finally, unlike prior gradient-free approaches that prune layer-by-layer (Dekhovich et al., 2021; Nova et al., 2023; Sun et al., 2023), Bonsai takes a holistic view across the entire model: modules across layers are removed and evaluated together, with relevance scores computed globally to make pruning decisions that better preserve accuracy.
Through extensive experiments, we demonstrate Bonsai’s effectiveness. When compared to other structured pruning methods that also avoid backward passes, including a forward-only structured pruning version of Wanda (Sun et al., 2023), Bonsai consistently produces models with better perplexity at any given speedup target. Wen the resulting pruned model small enough to be fine-tuned on available hardware, Bonsai compared to semi-structured pruning methods like Wanda (Sun et al., 2023)—which is also fine-tuned—Bonsai not only achieves similar performance, but also delivers 2×2$\times$2 × faster inference (an overall inference speed-up of 1.58×1.58$\times$1.58 × compared to the parent model). Even when compared to the state-of-the-art gradient-based structured pruning methods like LLM-Pruner and LoRAPrune, Bonsai produces models that outperform these more memory-intensive approaches on 4 out of 6 evaluations, despite using significantly less memory during pruning. Our results lead us to explore Bonsai’s practical utility by pruning the  3B Phi-2 (Li et al., 2023) model to a  1.8B model that performs competitively against other sub-2B parameter models on the Huggingface Open LLM leaderboard.
Overall, by reducing the memory and computational requirements of structured pruning, Bonsai enables more efficient adaptation of LLMs across a wider range of hardware configurations. This makes these LLMs more accessible while maintaining their performance advantages, representing an important step forward for practical LLM compression.
@@ 2 Methodology
We first cover background in LLM pruning, and then discuss our structured pruning method that exclusively performs inference on the parent model.
@@ 2.1 Background on Pruning, Problem Definition and Notation Setup
We are given an LLM, $\mathbf{M}_{\theta}$, parameterized by $\theta\in\mathbb{R}^{D}$. Also provided is $U$, a utility function that evaluates a model’s performance on a target task. We are interested in pruning $\mathbf{M}_{\theta}$ to produce a smaller and faster but still performant (with respect to $U$) model under the constraint that we only have enough memory to run inference on $\mathbf{M}_{\theta}$. Even though we assume we can run $\mathbf{M}_{\theta}$ on available hardware, pruning can be critical for achieving latency targets, reducing compute burden, or making the model small enough to adapt to new (out-of-domain) tasks by gradient-based fine-tuning.
Unstructured pruning approaches compress $\mathbf{M}_{\theta}$ by removing individual parameters $\theta_{j}$ from the model. This results in the updated model consisting of sparsified weight matrices with a smaller memory footprint. Unfortunately, the updated model does not enjoy inference speedups except when specialized hardware is available and thus poses a compute burden during inference. Whilst semi-structured variants – those that remove parameters in patterns like 2:4 or 4:8 (Mishra et al., 2021) – achieve some speedup, these are modest compared to those achieved with structured pruning.
Structured pruning takes a more modular view of the units to be removed from $\mathbf{M}_{\theta}$. Take that $\mathbf{M}_{\theta}$ is made up of modules $\mathbf{m}=\{m_{i}\}_{i\in[N]}$ each with corresponding parameter counts $\mathbf{s}=\{s_{i}\}_{i\in[N]}$ such that $\sum_{i}s_{i}=D$. For a transformer model, $\mathbf{m}$ could consist of attention heads, dimensions of fully connected layers, or even whole layers. For simplicity, we assume that $\mathbf{m}$ is made up of non-overlapping modules. Structured pruning compresses $\mathbf{M}_{\theta}$ by finding accurate sub-models defined by subsets of $\mathbf{m}$: provided with $\bar{\mathbf{m}}\subseteq\mathbf{m}$, we can construct an updated model $\mathbf{M}_{|\bar{\mathbf{m}}}$ that is produced by dropping the modules not in $\bar{\mathbf{m}}$ from $\mathbf{M}$. Thus, given a sparsity target $p$, structured pruning can be cast as the following combinatorial optimization problem:
@@
$\mathcal{F}_{p}$ consists of all sub-models that meet the sparsity threshold. Note that, in general, not only does $\mathbf{M}_{|\mathbf{m}^{\ast}}$ have a smaller memory footprint than $\mathbf{M}$, it is also faster to run inference on it since it has fewer modules. Many structured pruning methods attempt to solve Equation 1 by gradient-guided optimization (or search) over the space of sub-models. However, since we are interested in the memory-constrained setting where computing gradients is not feasible, these methods cannot be used.
Figure 2: Overview of Bonsai. Perturbations to the original model are generated according to an intelligent prior. These perturbed models are evaluated on downstream task and the collected data is used to build a simple linear model of module importances. Bonsai iteratively drops modules of least importance until the target model size is reached.
@@ 2.2 Estimating module relevance with only forward passes
We consider pruning a model too large to fit in memory, allowing only forward passes. Thus, we must solve Eq 1 using evaluations of $U$ rather than gradient-based optimization. A brute-force search over $\mathcal{F}_{p}$ is infeasible, as its size grows combinatorially with the model. For instance, pruning a single FC sub-layer in an LLM to 50% sparsity requires evaluating $\approx$ ($10^{4}\choose 10^{3}$ subsets.
We propose a computationally tractable approach where we first perform a small number, $n$, of evaluations, where $n\ll|\mathcal{F}_{p}|$, to gather data for estimating the relevance of each module in $\bm{M}$ with respect to the metric $U$. Upon carrying out this estimation, we can greedily choose the member of $\mathcal{F}_{p}$ that has the highest total module relevance. Specifically, let us assume that we have estimated $\bm{\beta}=\{\beta_{i}\}_{i\in[N]}$ to be the relevance of each of the $N$ modules. We can generate an approximate solution to Equation 1 as:
@@
$\begin{split}\hat{\bm{\beta}}&=\mathrm{argmin}_{\bm{\beta}\in\mathbb{R}^{N}}\quad\frac{1}{n}\sum_{\left(\bar{\bm{m}}_{k},U_{k}\right)\in\mathbb{D}}\left(U_{k}-\beta^{T}\alpha_{\bar{\bm{m}}_{k}}\right)^{2}+\gamma\|\bm{\beta}\|\\ \end{split}$ (3)

### v4

@@ Abstract
Structured pruning is a promising approach to create smaller, faster large language models. However, existing methods typically rely on computing the gradient via backward passes, which can inflate memory requirements and compute costs. In this work we introduce Bonsai, a gradient-free structured pruning method that eliminates the need for backpropagation, significantly reducing memory requirements and compute costs while achieving state-of-the-art pruning performance. Bonsai uses forward-pass-only perturbative pruning to enable efficient compression of large models on a broader range of hardware configurations. Unlike existing structured pruning approaches, Bonsai not only achieves better compression with fewer resources, but also produces models that are twice as fast as those generated by semi-structured pruning. As a concrete demonstration, we use Bonsai to prune 7B and 8B models to 50% sparsity on a single A6000 GPU—a task challenging for backprop-based methods in memory-constrained settings, as they require $2$-$3\times$ the memory. Our results show that removing backprop as a requirement not only enables pruning larger models on constrained hardware but can also lead to state-of-the-art efficiency and performance.
@@ 1 Introduction
Figure 1: Memory requirements can make pruning methods inaccessible. Gradient-based structured pruning methods require substantially more memory relative to forward-pass-only approaches like Bonsai, which can operate in memory-constrained settings (e.g., $\leq$24GB) while achieving superior performance among accessible methods. Results for LLaMA-2 7B at 50% sparsity on Wikitext-2.
The increasing scale of large language models (LLMs) has amplified the computational and memory requirements for their deployment and adaptation, presenting a significant barrier to widespread adoption. While structured pruning has emergeding methods require substantially more memory relative to as a popular approach to create smaller, faster models (Wang et al., 2019; Xia et al., 2022a), many existing methods rely on backpropagation, which significantly inflates memory and compute costs: backward passes consume $\gtrapprox 2\times$ (Bridger, 2023) the memory of a forward pass, with popular stateful optimizers like AdamW (Loshchilov and Hutter, 2017) requiring $\gtrapprox 3\times$ more memory. This often makes them impractical for a wide range of practitioners who operate under resource constraints, such as students, researchers, and small organizations with limited access to enterprise-grade, multi-GPU systems.
In response to this challenge, we present Bonsai, a structured pruning approach that operates exclusively with forward passes through the parent model. By eliminating the need for backward passes for pruning, Bonsai dramatically reduces memory requirements to enable the compression of models on commodity hardware and in memory-constrained environments—potentially obviating the need for more expensive, enterprise-grade solutions. Notably, Bonsai maintains lower memory requirements while still outperforming existing structured and unstructured pruning approaches in terms of compression/accuracy trade-offs (Fig 1).
The key challenge in structured pruning is determining which modules (e.g., attention heads, layer dimensions) are most important for model performance and can be safely removed. Rather than computing gradients to determine module importance, Bonsai employs an innovative perturbative approach, estimating module importances by evaluating the performance of sub-models using only inference.
We make this perturbative approach tractable through several key innovations. First, we formulate module importance estimation as an underdetermined regression problem, allowing us to infer the importance of a large number of modules by exploring a manageable number of random sub-models. This fundamentally differs from previous perturbative approaches, which require evaluating approximately as many sub-models as there are modules (Ancona et al., 2020), making them intractable for LLMs. Second, we use informative priors derived from existing pruning approaches (Han et al., 2015; Sun et al., 2023; An et al., 2024) to guide sub-model exploration, yielding better estimates of module relevance with fewer evaluations, instead of instantiating sub-models by dropping modules with equal likelihood (Kang et al., 2023). Finally, unlike prior approaches that prune layer-by-layer (Dekhovich et al., 2021; Nova et al., 2023; Sun et al., 2023; Ling et al., 2024), Bonsai takes a holistic view across the entire model: modules across layers are removed and evaluated together, with relevance scores computed globally to make pruning decisions that better preserve accuracy than the sub-optimality of localized methods.
Through extensive experimentation, we demonstrate that Bonsai’s approach is highly efficient and effective. When compared to other structured pruning methods that also avoid backward passes, including FLAP (An et al., 2024) and a forward-only structured pruning version of Wanda (Sun et al., 2023), Bonsai consistently produces models with better perplexity at any given speedup target. Critically, even when compared to the state-of-the-art gradient-based structured pruning methods like LLM-Pruner and LoRAPrune, Bonsai produces models that majorly outperform these more memory-intensive approaches, despite using significantly less memory during pruning. Our results lead us to explore Bonsai’s practical utility by pruning the 3B Phi-2 (Li et al., 2023) model to a 1.8B model that performs competitively against other sub-2B parameter models on the Huggingface Open LLM leaderboard.
A key advantage of Bonsai is that it not only reduces memory requirements during pruning but also unlocks the opportunity for efficient subsequent post-pruning adaptation. As the pruned model is small enough to fit on the same hardware that was used for inference, practitioners can then perform parameter-efficient fine-tuning to further recover performance. This capability addresses a fundamental misconception that gradient-based methods are always superior: in many cases, Bonsai is a prerequisite for reaching a stage where such fine-tuning is even possible, expanding access to the entire model adaptation pipeline and yielding superior end performance. Our approach thus provides a tangible solution for practitioners who do not have access to the high-cost resources typically associated with LLM development.
2 Related Work
We first discuss relevant work in LLM pruning and other complementary compression methods.
2.1 Unstructured Pruning
While structured pruning removes entire components like layers (Xu et al., 2020; Xia et al., 2022b), dimensions of linear layers (Wang et al., 2019) or attention heads (Michel et al., 2019; Held and Yang, 2022), unstructured pruning (Han et al., 2015; Frankle and Carbin, 2018; Benbaki et al., 2023; Sun et al., 2023) removes individual parameters of the model. These approaches achieve memory savings by inducing sparsity in the model weights, but they generally do not result in tangible model speedups except when specialized hardware is available (Mishra et al., 2021). Proposed semi-structured sparsity methods (Mishra et al., 2021) such as 2:4 and 4:8 patterns can provide faster inference, but the speedup gains they achieve are far from the idealized $2\times$.
2.2 Gradient-Based Structured Pruning
Most existing structured pruning techniques for large (over 1B scale) language models rely on gradient computation to estimate module importance. LLM-Pruner (Ma et al., 2023) uses first-order Taylor expansion to measure the importance of coupled structures in transformers, followed by LoRA fine-tuning for performance recovery. LoRAPrune (Zhang et al., 2024) reduces memory requirements compared to LLM-Pruner by using LoRA gradients rather than full model gradients for importance estimation, though it still requires gradient computation during the pruning process. Sheared LlaMA (Xia et al., 2024) takes a different approach by learning pruning masks during continued pre-training, integrating structured pruning with the training process. More recently, SparseLLM (Bai et al., 2024) decomposes the global pruning problem into manageable subproblems using auxiliary variables, though it maintains reliance on gradient information.
Ultimately, a fundamental limitation of gradient-based approaches is their memory requirements. Computing gradients for large models during pruning significantly increases memory usage beyond what is needed for inference alone, making these methods challenging to apply on resource-constrained hardware. In this work we show that it is possible to remove these additional memory requirements while still achieving comparable or superior performance to gradient-based pruning methods.
2.3 Memory-Efficient Structured Pruning Approaches
The memory constraints of gradient-based methods have motivated research into gradient-free alternatives, though such work remains more limited for LLMs. SliceGPT (Ashkboos et al., 2024) utilizes PCA to remove up to 25% of weight matrices, although it selects embedding reduction directions based on variance, which may not correlate with actual model utility for downstream tasks. In the same vein, SlimGPT (Ling et al., 2024) extends the classical Optimal Brain Surgeon (OBS) framework (Hassibi and Stork, 1992) to LLMs to reduce memory costs; however, the work admits that its layer-wise nature can lead to ‘error accumulation’ and an inability to use global information when pruning; we find that this makes it substantially less performant than Bonsai’s global approach.
For smaller language models like BERT, Nova et al. (2023) proposed Kernelized Convex Masking (KCM) for gradient-free structured pruning. Unfortunately, to prune a fully connected layer with $K$ intermediate units, KCM requires instantiating a $K\times K$ coefficient matrix for each layer. While this is feasible for BERT models, a typical LLM (AI, 2023; Touvron et al., 2023) has $K\sim 10^{4}-10^{5}$ which would make the size of each per-layer coefficient matrix comparable to the size of the LLM itself. In computer vision, there exist several perturbative gradient-free structured pruning techniques (Ancona et al., 2020; Dekhovich et al., 2021). However, these methods have been exclusively applied to small vision models (e.g., VGG (Simonyan and Zisserman, 2014) and WideResnet (Zagoruyko and Komodakis, 2016)) and would not scale to LLMs as-is.
Most comparably, FLAP (An et al., 2024) introduces a fluctuation-based metric that operates without gradients using statistical fluctuations in activations to estimate module importance. We show in Section 4.4 and Appendix E.3 that Bonsai’s flexible framework outperforms FLAP and can in fact incorporate FLAP’s fluctuation-based importance metric as a special case, being a more general methodology. Finally, Probe Pruning (Le et al., 2025) introduces a different approach with “online, dynamic, structured pruning" that operates at inference time via model probing, and LoRAM (Zhang et al., 2025) trains LoRA adapters on a pruned model and uses recovered low-rank matrices with the original model for inference. Both of these methods tackle different problems to Bonsai—online pruning/fine-tuning rather than one-time permanent pruning—but they demonstrate the broader, shared trend toward memory-efficient model adaptation.
2.4 Positioning of Bonsai
Overall, the existing literature shows a clear gap: there is a need for methods that can operate within memory-constrained environments while achieving competitive structured pruning performance for LLMs. The highest-performing structured pruning approaches require computational resources that exceed what is available on single consumer GPUs.
Our approach addresses this gap by formulating structured pruning as an underdetermined regression problem that can be solved using only forward passes through the model. Key distinctions of our approach include:
  • Unlike methods requiring gradient computation or auxiliary mathematical operations like Hessian storage, Bonsai operates solely through model inference.
  • Rather than making layer-wise pruning decisions, our method estimates module importance globally across the entire model through perturbative evaluation.
  • Our approach requires only the memory needed for inference plus a small overhead for storing perturbation results, making it accessible on standard hardware.
  • The regression-based formulation provides theoretical grounding for estimating module importance from limited samples, avoiding purely heuristic approaches.
2.5 Additional Methods for LLM Compression
Finally, although we focus on structured pruning in this work, we note that prior research has explored various other compression schemes such as distillation (Hinton et al., 2015; Gu et al., 2023) and quantization (Xiao et al., 2023) to create smaller models from larger pre-trained ones. Similar to structured pruning, these compression methods themselves often impose significant computational burdens. For example, distillation-based techniques require using LLMs to generate large amounts of teacher data (Jiao et al., 2019; Hsieh et al., 2023). Although unstructured pruning (Frantar and Alistarh, 2023; Sun et al., 2023) and quantization have lower training-time resource demands, the models they produce either require specialized hardware to achieve speedups (Mishra et al., 2021) or may actually slow down inference due to additional computational overhead (Dettmers et al., 2022). Table 1 summarizes these characteristics across different compression approaches.
@@ 2.5 Additional Methods for LLM Compression
Regime Resource Approaches
Train Memory
Compute
Inference Memory
Compute
Overall, each approach has distinct hardware requirements and performance characteristics that make them suitable for different deployment scenarios. Our structured pruning approach could potentially be combined with these techniques; for instance, quantization could be applied post-pruning, and structured pruning could serve as initialization for distillation approaches. We leave such combinations for future work and focus specifically on the forward-pass-only structured pruning problem.
@@ 3 Methodology
We cover background in LLM pruning (§3.1), and then discuss Bonsai, our structured pruning method that exclusively performs inference on the parent model to prune. Figure 2 provides an overview of our approach as detailed in the following subsections: §3.2 describes our perturbative estimation; §3.3 explains our informative prior-based sampling strategy; and §3.4 discusses our iterative pruning procedure.
@@ 3.1 Background on Pruning, Problem Definition and Notation Setup
We assume that we are given an LLM, $\mathbf{M}_{\theta}$, parameterized by $\theta\in\mathbb{R}^{D}$. Also provided is $U$, a utility function that evaluates the model’s performance on a target task. We instantiate $U$ as language modeling perplexity: Wikitext-2 training for Wikitext-2 experiments and C4 for zero-shot tasks. We are interested in pruning $\mathbf{M}_{\theta}$ to produce a smaller and faster but still performant (with respect to $U$) model under the constraint that we only have enough memory to run inference on $\mathbf{M}_{\theta}$. Even though we assume we can run $\mathbf{M}_{\theta}$ on available hardware, pruning can be critical for achieving latency targets, reducing compute burden, or making the model small enough to adapt to new (out-of-domain) tasks by performing gradient-based fine-tuning.
Unstructured pruning approaches compress $\mathbf{M}_{\theta}$ by removing individual parameters $\theta_{j}$ from the model. This results in the updated model consisting of sparsified weight matrices with a smaller memory footprint. Unfortunately, the updated model may not enjoy inference speedups except when specialized hardware is available and thus can pose a compute burden during inference Mishra et al. (2021). While semi-structured variants – those that remove parameters in patterns like 2:4 or 4:8 (Mishra et al., 2021) – achieve some speedup, these are modest compared to those achieved with structured pruning.
Structured pruning takes a more modular view of the units to be removed from $\mathbf{M}_{\theta}$. Consider that $\mathbf{M}_{\theta}$ is made up of modules $\mathbf{m}=\{m_{i}\}_{i\in[N]}$ each with corresponding parameter counts $\mathbf{s}=\{s_{i}\}_{i\in[N]}$ such that $\sum_{i}s_{i}=D$. For a transformer model, $\mathbf{m}$ could consist of attention heads, dimensions of fully connected layers, or even whole layers. For simplicity, we assume that $\mathbf{m}$ is made up of non-overlapping modules. Structured pruning compresses $\mathbf{M}_{\theta}$ by finding accurate sub-models defined by subsets of $\mathbf{m}$: provided with $\bar{\mathbf{m}}\subseteq\mathbf{m}$, we can construct an updated model $\mathbf{M}_{|\bar{\mathbf{m}}}$ that is produced by dropping the modules not in $\bar{\mathbf{m}}$ from $\mathbf{M}$. Thus, given a sparsity target $p$, structured pruning can be cast as the following combinatorial optimization problem:
@@
$\mathcal{F}_{p}$ consists of all sub-models that meet the sparsity threshold. Note that, in general, not only does $\mathbf{M}_{|\mathbf{m}^{\ast}}$ have a smaller memory footprint than $\mathbf{M}$, it is also faster to run inference on it since it has fewer modules. Many structured pruning methods attempt to solve Equation 1 by gradient-guided optimization (or search) over the space of sub-models. In this work, we instead explore approaches suitable for memory-constrained settings, where computing gradients may be expensive or infeasible.
Figure 2: Bonsai estimates module importance through regression on perturbative evaluations. Rather than requiring as many sub-models as there are modules (intractable for LLMs), we solve an underdetermined regression problem: sample $n\ll N$ sub-models according to informative priors, evaluate each via for forward pass, then regress to estimate global importance scores $\bm{\beta}$. Black squares indicate removed modules.
@@ 3.2 Estimating module relevance with only forward passes
We specifically aim to develop an approach for pruning that only relies on computing the forward pass over the model. Thus, we must solve Eq 1 using evaluations of $U$ rather than gradient-based optimization. A brute-force search over $\mathcal{F}_{p}$ is infeasible, as its size grows combinatorially with the model. For instance, pruning a single FC sub-layer in an LLM to 50% sparsity requires evaluating $\approx$ $10^{4}\choose 10^{3}$ subsets.
We instead propose a computationally tractable approach where we first perform a small number, $n$, of evaluations, where $n\ll|\mathcal{F}_{p}|$, to gather data for estimating the relevance of each module in $\bm{M}$ with respect to the metric $U$. Upon carrying out this estimation, we can greedily choose the member of $\mathcal{F}_{p}$ that has the highest total module relevance. Specifically, let us assume that we have estimated $\bm{\beta}=\{\beta_{i}\}_{i\in[N]}$ to be the relevance of each of the $N$ modules. We can generate an approximate solution to Equation 1 as:
@@
[…]
@@
Algorithm 1 Bonsai Pruning Method
1:Input:
2:Model $\mathbf{M}_{\theta}$], sub-models per iteration $n_{\mathrm{iter}}$]
3:Sparsity per iteration $p_{\mathrm{iter}}$], Target sparsity [$p$]
4:Module list $\mathbf{m}$]
5:
6:for $l=1$ to $\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ do
7:  $\mathbb{\rho}^{l}\leftarrow$ Calculate unstructured pruning metric for all modules in $\mathbf{m}$
8:  $\bar{\mathbb{\rho}}^{l}\leftarrow$ Fix the top $(1-2p_{\mathrm{iter}})$ of $\mathbb{\rho}^{l}$ to $\infty$
9:  Sample $\{\mathbf{\bar{m}}_{i}\}_{[n_{\mathrm{iter}}]}$ sub-models according to $\bar{\mathbb{\rho}}^{l}$
10:  Run forward pass on each sub-model and compute $U$. Construct $\mathbb{D}^{l}=\left\{\bar{\mathbf{m}}_{i},U_{i}\right\}_{[n_{\mathrm{iter}}]}$
11:  $\beta^{\l }\leftarrow\mathrm{Regress}\left(\mathbb{D}^{l}\right)$
12:  $\{m_{\mathrm{pruned}}\}\leftarrow$ sort $\beta^{\l }$ and drop the bottom $k$ modules that make up $p_{\mathrm{iter}}$ fraction of the model.
13:  $\mathbf{m}\leftarrow$ update module list to exclude $\{m_{\mathrm{pruned}}\}$
14:end for
15:Output: Pruned model $\mathbf{M}_{|\mathbf{m}}$